A closed system has 150 KJ of work supplied when the initial volume is 0.6m^3 and pressure of the system changes in the relation p = 8-4V, where V is in m^3 and p is in bar. Compute the final pressure and volume of the system. Express your answer in Pa and m^3


1
Expert's answer
2021-01-18T01:18:46-0500

W = nRT ln (p0/pf) = pV ln (p0/pf)

p = 8 - 4V = 8 - 4 x 0.6 = 5.6 bar

150 x 103 J = 5.6 x 105 Pa x 0.6 m3 x ln (5.6 x 105 / pf)

pf = 3.6 x 105 Pa

3.6 = 8 - 4Vf

Vf = 1.1 m3